OpenStudy (anonymous):

Use the proportion x/z=6/5. COmplete each statement, Justify your answer. 10.) x/6= 11.) x+z/z= 12.)z/x= 13.) 5x= THIS IS SO CONFUSING TO ME!!! PLEASE HELP

OpenStudy (mathstudent55):

Here are some rules of proportions: If a/b = c/d is true, then a/c = b/d, b/a = d/c, d/b = c/a are also true If a/b = c/d, then (a + b)/b = (c + d)/d

OpenStudy (anonymous):

if x/z = 6/5 then x=6 and z=5 right?

OpenStudy (anonymous):

what? I'm going to cry... I'm so lost

OpenStudy (anonymous):

look at what the question says

OpenStudy (anonymous):

x=6 so plug it in

OpenStudy (anonymous):

so for one you put it as 6/6 because you plugged 6 in for x?

OpenStudy (anonymous):

remember x=6 z=5 10. 6/6=1 11.(6+5)/5 =2.2 12.5/6= 0.83 13.5(6) or 5*6=30

OpenStudy (mathstudent55):



OpenStudy (anonymous):

but it wants another proportion

OpenStudy (mathstudent55):

If you look at the proportion written with fractions as I did above in the drawing it'll be easier to understand.

OpenStudy (mathstudent55):

The first rule of poroprtions I wrote above shows that a proportion can be read in any direction.

OpenStudy (anonymous):

ok then you would just use my fractions

OpenStudy (anonymous):

Okay I got it thanks

OpenStudy (mathstudent55):